### v2

# How Should I Build A Benchmark? Revisiting Code-Related Benchmarks For LLMs

Abstract
Various benchmarks have been proposed to assess the performance of large language models (LLMs) in different coding scenarios. We refer to them as code-related benchmarks. However, there are no systematic guidelines by which such a benchmark should be developed to assure its quality, reliability, and reproducibility. We propose How2Bench comprising a 55-criteria checklist as a set of guidelines to comprehensively govern the development of code-related benchmarks. Using How2Bench, we profiled 274 benchmarks released within the past decade and found concerning issues. Nearly 70% of the benchmarks did not take measures for data quality assurance; over 10% did not even open source or only partially open source. Many highly cited benchmarks have loopholes, including duplicated samples, incorrect reference codes/tests/prompts, and unremoved sensitive/confidential information. Finally, we conducted a human study involving 49 participants and revealed significant gaps in awareness of the importance of data quality, reproducibility, and transparency. For ease of use, we provide a printable version of How2Bench in Appendix E.
[…]
MyFramelinecolor=black, outerlinewidth=0.3pt, roundcorner=5pt, skipabove = 5.5pt, skipbelow = 5.5pt, innertopmargin=5.5pt, innerbottommargin=5.5pt, innerrightmargin=5.5pt, innerleftmargin=5.5pt, backgroundcolor=bg,
How Should I Build A Benchmark?
Revisiting Code-Related Benchmarks For LLMs
Jialun Cao1 , Yuk-Kit Chan2thanks:   Both authors contributed equally to this research. , Zixuan Ling211footnotemark: 1 , Wenxuan Wang1thanks:   Wenxuan Wang is the corresponding author., Shuqing Li2, Mingwei Liu3, Ruixi Qiao4, Yuting Han5, Chaozheng Wang2, Boxi Yu6, Pinjia He6, Shuai Wang1, Zibin Zheng3, Michael R. Lyu2 , Shing-Chi Cheung1 1The Hong Kong University of Science and Technology 2The Chinese University of Hong Kong 3Sun Yat-Sen University 4 Chinese Academy of Science, Institute of Automation 5 Beijing Language and Culture University 6The Chinese University of Hong Kong, Shenzhen jcaoap@cse.ust.hk, jwxwang@gmail.com
1 Introduction
If you cannot measure it, you cannot improve it. ❞ — Lord Kelvin (1824-1907)
Recent large language models (LLMs) have shown remarkable capabilities across various domains such as software development Chen et al. (2021a), question answering Rogers et al. (2023), and math reasoning (Imani et al., 2023). Various benchmarks Chen et al. (2021a); Jimenez et al. (2024); Austin et al. (2021); Yue et al. (2024); Du et al. (2023a) are proposed to evaluate LLMs’ effectiveness and limitations from multiple perspectives in different application scenarios.
However, doubts regarding the quality, reliability, and transparency of various code-related benchmarks arise. For example, a recent study pointed out that “current programming benchmarks are inadequate for assessing the actual correctness of LLM-generated code” (Liu et al., 2023a). Other accusations, including irreproducible (Reuel et al., 2024), closed data sources (Cao et al., 2024b), low quality Qiu et al. (2024a); Yadav et al. (2024a), and inadequate validation measures (Liu et al., 2023a), were also raised, undermining the credibility of these benchmarks and thereby their subsequent evaluation results. This motivates the need for rigorous and thorough guidelines to govern code-related benchmark development.
In this paper, we introduce How2Bench, a comprehensive guideline consisting of a 55-criteria checklist specially designed for code-related benchmarks. This checklist covers the entire lifecycle of benchmark development, from design and construction to evaluation, analysis, and release as shown in Figure 1. It underwent multiple iterations — we initiated a draft inspired by open-source software guidelines Fogel (2005) and classical measurement theory Suppes et al. (1962). We refined it through iterative discussions with practitioners, leading to the finalization of these criteria. How2Bench emphasizes reliability, validity, open access, and reproducibility in the benchmark development, ensuring high standards and fostering a more reliable and transparent environment.
Following How2Bench, we conducted an in-depth profiling of 270+ code-related benchmarks developed over the past decade (2014 - 2024). The extent of criteria violations by the profiled benchmarks is concerning:
⚫ Almost 70% of the benchmarks did not take any measures for data quality assurance;
⚫ Over 90% did not consider code coverage when use passing test cases as an oracle;
⚫ Over half of the benchmarks did not provide the essential information (e.g., experiment setup, prompts) for reproducibility;
⚫ Over 10% are not open source or only partially open source.
We observed that even highly cited benchmarks have loopholes, including duplicated samples, incorrect reference/tests, unclear displays, and unremoved sensitive/confidential information. We also observed these loopholes can propagate. Over 18% of the benchmarks serve as data sources (Figure 8) for subsequent benchmarks. Therefore, the data quality of benchmarks affects their credibility and likely impacts future benchmarks.
To understand the usefulness of the criteria in How2Bench, we conducted a human study involving 49 participants through questionnaires. All participants concurred on the necessity of having a checklist for benchmark construction to enhance quality. Nearly all participants with experience in benchmark development acknowledged the importance of all these 55 criteria. The study also exposed gaps in quality awareness: 16% of participants were unaware of the necessity for data denoising, and over 40% were not aware that the experimental setup and environment could impact the reproducibility and transparency. This paper makes contributions in five aspects:
  • Novelty. We introduce How2Bench, a comprehensive set of guidelines packaged as a 55-criteria checklist that covers the lifecycle of code-related benchmark development.
  • Significance. How2Bench presents the first comprehensive set of actionable guidelines for developing high-quality benchmarks, striving to create a more reliable and transparent environment. The human study also highlighted the demand for such a detailed guideline.
  • Usefulness. How2Bench serves as a guideline for practitioners before/during developing code-related benchmarks, and a checklist for evaluating existing benchmarks after their release. For ease of use, we also provide a printable version of How2Bench on Appendix E.
  • Generalizability. Most criteria listed in How2Bench can be adopted or adapted to other benchmarks such as Question-answering, mathematical reasoning, and multi-modal benchmarks.
  • Long-term Impact. Our statistics alert the community to the severity and prevalence of non-standard practices in benchmark development. It ultimately improves the overall quality of benchmarks due to the propagation effect among them.
2 Background
2.1 Code-related Benchmarks
Benchmarks for coding tasks like code generation Chen et al. (2021a); Austin et al. (2021), defect detection Just et al. (2014); Gao et al. (2023b); Liu et al. (2024c), and program repair Jimenez et al. (2024); Risse and Böhme (2024) are increasingly common, reflecting the growing needs for using LLMs for coding tasks. Recent studies have highlighted various issues with these benchmarks, ranging from design inconsistencies to scope and applicability limitations. For example,  (Liu et al., 2023a) found that even some widely used benchmarks, such as HumanEval (Chen et al., 2021a) and MBPP Austin et al. (2021), contains a non-trivial proportion of bugs in implementation, documentation, and test cases. Our work, in comparison, introduces a detailed guideline that guides the benchmark development during the entire lifecycle.
Figure 1: Lifecycle of Benchmark Development
2.2 Related Studies and Surveys
Several recent surveys and empirical studies have profiled the status quo of LLM development. These studies either explore the overall performance for certain areas such as software engineering Hou et al. (2023); Wang et al. (2024a) or investigate the capabilities of LLMs on specific tasks such as code generation Dou et al. (2024); Yu et al. (2024) and test generation Schäfer et al. (2024); Yuan et al. (2024, 2023a). A survey Chang et al. (2024) about how to evaluate LLMs was proposed to answer what/where/how to evaluate LLMs. This paper differs from these studies in its purpose and perspectives. Unlike these benchmarks, our work proposed guidelines for future benchmark development and provided a checklist to assess the quality of these existing benchmarks.
Recently, BetterBench Reuel et al. (2024) is a concurrent work assessing the AI benchmarks against 46 criteria. Then, it scored 24 AI benchmarks in various domains and ranked them. BetterBench differs from this paper in several key aspects: scope (general benchmarks vs. code-related benchmarks), lifecycle division (it addresses benchmark retirement, while How2Bench focuses on benchmark evaluation, analysis, and release), and objectives (scoring benchmarks vs. offering comprehensive guidelines for future benchmark development). Additionally, the study in this paper was conducted on a much larger scale (24 vs. 274 benchmarks), statistically highlighting the prevalent issues in existing benchmarks.
3 Design
3.1 The Lifecycle of Benchmark Development
Code-related benchmark development comprises five typical phases (Phase 0 - 4), as shown in Figure 1, explained in detail as follows.
Phase 0. Design. At the beginning of benchmark development, it is vital to identify the motivation, the scope and the capabilities required by the application scenario of interest.
Phase 1. Construction. After establishing the motivation and purpose, the Benchmark Construction phase moves from design to execution. Typically, data is collected from public coding websites such as GitHub, LeetCode and StackOverflow. This is followed by preprocessing, which includes filtering, cleaning (e.g., deduplication, denoising), and curation (e.g., aligning tests with corresponding code). The phase usually ends with a validation process, which can be manual or automated.
Phase 2. Evaluation. Once the benchmark is available, the next step is to apply it to LLMs, validating if it can effectively measure the intended LLM capabilities. Essential considering factors include selecting a representative array of LLMs, configuring settings like prompts and hyperparameters for consistency, choosing appropriate experimental environments to meet LLM requirements, and implementing thorough logging to ensure dependable and reproducible results.
Phase 3. Analysis. After evaluation, experimental results are analyzed, drawing conclusions on LLMs’ capabilities. This phase involves comparing each LLM’s performance to identify standout or underperforming models. Then, proper visual aids such as bar charts and tables can be used to display the experimental results, presenting clearer observation and deeper inspiration, such as the correlations between models, the correlations with related benchmarks, or performance in upper-/down-stream tasks. Indeed, a thorough analysis helps pinpoint areas for improvement and guides future enhancements in LLM development.
Phase 4. Release. The final phase is to make the benchmark open-accessible. This phase involves meticulously preparing all materials associated with the benchmark, ensuring they are ready for open access to foster widespread adoption and collaboration. Clear, comprehensive documentation is provided to guide users on effectively utilizing the benchmark. Additionally, all logged experiment details are made available, enhancing the reproducibility and transparency of the benchmark.
Figure 2: Workflow of study process
3.2 Study Design
Our study consists of four steps (Figure 2).
Step 1. Guideline Construction. To begin with, we sketched the initial guidelines for each phase in the benchmark development lifecycle (Section 3.1, Figure 1) by reviewing existing literature (Suppes et al., 1962; Zheng et al., 2023b; Schäfer et al., 2024; Reuel et al., 2024) and brainstorming. After that, we refined the guidelines through a series of interviews with various stakeholders, including model developers and benchmark builders, allowing for the addition, deletion, or modification of criteria based on expert feedback and practical insights. This phase concludes with the finalization of our guidelines, How2Bench. This detailed checklist consists of 55 criteria over the benchmark lifecycle, providing effective guidelines for rigorous and reliable benchmark development.
Step 2. Literature Profiling. This step begins by collecting related benchmarks according to their publication time, venue, and coding tasks, and then employing techniques like snowballing to ensure a comprehensive collection. This step leads to 274 code-related benchmarks for study. The detailed statistics can be found in the Appendix D. This step is followed by profiling each selected benchmark through a thorough review of corresponding papers and examination of the released artifacts or homepages associated with these benchmarks. The phase is completed by reporting statistics that highlight overall trends, advantages, and disadvantages identified during the profiling, providing a clear and structured overview of existing benchmarks.
Step 3. Focused Case Study. After obtaining an overall impression of existing benchmarks, we selected 30 (= 5 * 6) representative benchmarks from top-5 tasks, with top-5 highly-cited benchmarks plus the latest 1 benchmark (Appendix C). Each selected benchmark is then analyzed against How2Bench, examining how well they meet the established criteria, studying their overall statistics, and identifying both exemplary and poor cases. Insights and references from existing literature are also incorporated to enrich the analysis, providing a deeper understanding of the benchmarks’ performance and areas for improvement.
Step 4. Human Study. The final step is a human study that evaluates the importance and practicality of How2Bench. This involves designing a questionnaire to gather diverse, logical insights, which is then distributed to a targeted audience. After collecting and filtering responses for quality, the data is analyzed to derive insights. See Appendix B for details.
4 Guideline – “How2Bench
The completed guideline How2Bench with 55 criteria can be found in Appendix E.
Figure 3: Guideline for Benchmark Design
4.1 Guideline for Benchmark Design
Explanation – For benchmark design, we listed four essential criteria, as shown in Figure 3. In particular, the guideline starts by recommending that benchmarks should initially assess if they are addressing a significant gap in existing research, ensuring the relevance and necessity of the benchmark. The scope of the benchmark is expected to be well-defined, clarifying the capabilities or characteristics being tested, how these relate to practical scenarios such as programming assistance or automated testing, and the relevance of these capabilities in real-world applications.
Key Statistics – According to our statistics among 270+ benchmarks, apparent research bias can be observed in terms of coding tasks, programming languages, and code granularities are observed (Appendix A.1). For example, as shown in Figure 11, 36.13% (99/274) are code generation benchmarks, followed by program repair 9.85% (27/274).
Also, during the focused case study (listed in Appendix C), we identified that 10% benchmarks have not explicitly specified the capabilities (e.g., intention understanding, program synthesis) to be evaluated, and 30% have not specified application scenarios the benchmark targets.
Besides, we also identified a case in MBPP Austin et al. (2021) where a case fell out of the target evaluation capabilities (Appendix A.2). Indeed, clearly defining the application scenarios/scopes/capabilities could help benchmark constructors establish precise goals for the design and development of the benchmark, ensuring accuracy in the evaluation.
Lastly, Figure 12 shows that 58% (158/274) code-related benchmarks involve Python, followed by 39% (107/274) involving Java. In comparison, 31 programming languages are only covered by one benchmark, and less than five benchmarks cover other 19 programming languages. This observation consolidates the observation from previous works Cao et al. (2024a); Hou et al. (2023) on a larger scale.
{mdframed}
[style=MyFrame] ▲ Severity – Current benchmarks exhibit an apparent imbalance in coding tasks and programming languages dominated by code generation and Python, leaving research blanks to be filled. Also, even highly cited benchmarks may have samples that do not fall into the examined capabilities.
4.2 Guideline for Construction
Figure 4: Guideline for Benchmark Construction
Explanation – Figure 4 shows 19 criteria for benchmark construction. Essentially, for data source, the key considerations include verifying the traceability and quality of the data source, addressing potential data contamination Sainz et al. (2023), and ensuring that the data sampling processes are scientifically robust and rigorous. Also, for data representativeness, it also guides through specific checks to ensure the benchmark’s scope is strictly adhered to, such as making sure every data point falls within the targeted scope and that the data can cover all studied capabilities, domain knowledge, and application scenarios.
For data preprocess and cleaning, it also stresses handling code-specific aspects, such as compilability and execution, along with cleaning and manually reviewing data for quality assurance. Output validation methods and evaluation metrics must be carefully designed and reviewed to ensure they effectively measure the benchmark’s goals. Lastly, it suggests considering additional evaluation perspectives, such as safety and security checks, ensuring the code does not contain sensitive information.
Key Statistics – According to our statistics (Appendix A.3), the 270+ benchmarks exhibit numerous irregularities in their implementation, which could significantly threaten the reliability of the benchmarks. Surprisingly, 62% of benchmarks did not deduplicate or did not mention. Near 80% benchmarks did not consider or handle data contamination threats. About 70% of the benchmarks did not go through any quality assurance checks such as manual checks and code execution. In particular, we summarized the commonly-used data quality assurance metrics and their frequency: manual check (22.6%), code execution (2.2%), LLM check (1.5%), others (e.g. the number of stars or heuristic rules, 5.8%).
Also, since we focus on code-related benchmarks, which usually accompany test cases, test coverage also needs to be considered. As pointed out by prior study Liu et al. (2023a), inadequate test coverage can lead to inflated evaluation results. However, according to our statistics, we observed that only 8.7% of benchmarks have considered test coverage when using test cases as oracles. It severely affects the reliability of findings on these benchmarks, potentially misguiding future research and application decisions based on these flawed assessments.
{mdframed}
[style=MyFrame] ▲ Severity – Most benchmarks display severe loopholes in data preparation and curation processes. Essential steps such as data cleaning, process oversight, and quality checks are often neglected.
4.3 Guideline for Evaluation
Figure 5: Guideline for Benchmark Evaluation
Explanation – Guidelines for benchmark evaluation focus on the rigorousness and reliability of the evaluation. How2Bench provides 12 criteria for benchmark evaluation, as shown in Figure 5. It mainly focuses on the comprehensive evaluation processes for benchmarks involving LLMs. For evaluation design, it stresses the importance of assessing a sufficient and representative range of LLMs to ensure the benchmark’s applicability across various model families and configurations, both open and closed-source. Figure 29 and Figure 30 shows the distribution of numbers of LLMs studied and the most exercised LLMs.
Also, prompting has a direct impact on the quality of the LLMs’ output results Wei et al. (2022); He et al. (2024a); Jin et al. (2024); Ye et al. (2023). As pointed out by a recent study, up to 40% performance gap could be observed in code translation when prompts vary He et al. (2024b).
Additionally, the experiment environment is essential for reproducibility and transparency. Indeed, the hardware, software, and platform environments used during experiments might influence the outcomes Ghosh (2024). Furthermore, because of the nondeterministic nature of LLMs, experiments should be repeated, and randomization strategies should be used to mitigate the effects of randomness and parameter configuration biases. Lastly, meticulously documented logs of the experimental process are advised to facilitate transparency and reproducibility, detailing everything from parameter settings to the specific LLM frameworks used.
Key Statistics – Among the 274 benchmarks, 183 of them are evaluated over LLMs. According to our statistics (Figure 29), over 34% of the benchmarks were evaluated on fewer than 3 LLMs, with 11.48% benchmarks only evaluated on one LLM. Such evaluation results can hardly be generalized to other LLMs. Furthermore, more than half of the benchmarks studied fewer than 6 LLMs (51% = (21 +22 + 20 + 4 + 12+15)/183).
☛ For reference, we listed the top 10 most studied LLM families in Figure 30. Among them, the GPT and CodeLlama series are the most extensively studied, accounting for 63% (116/183) and 33% (60/183), respectively. Under the constraints of time and available resources, it is beneficial to evaluate more representative LLMs.
The prompt quality also greatly impacts the LLM evaluation He et al. (2024b). According to a recent study, up to 40% performance vary could be observed in code translation task He et al. (2024b). So, carefully designing a prompt needs consideration. However, 73.3% representative benchmarks (Appendix C) do not validate whether the prompts they used are well-designed (Appendix A.4). Similarly, though 94.9% benchmarks were evaluated in a zero-shot manner, only 21.2% benchmarks were evaluated under few-shot, 8.8% under Chain-of-Thought and 2.6% under RAG (Appendix A.4).
Regarding the evaluation process, our statistics exposed that only 35.4% of benchmark evaluations have been repeated (Appendix A.4). Also, regarding the transparency and matriculated documents, the observation is not optimistic – Only 3.6% benchmarks provided their experiment environment. More than 50% of benchmarks did not provide reproducible instructions such as prompts, examples for few-shot learning, or content for retrieval (Figure 39). Less than half (42.7%) provide hyperparameters such as temperature for reproduction.
{mdframed}
[style=MyFrame] ▲ Severity – Over 60% of evaluations have not been repeated to eliminate the impact of randomness. Only a few (less than 3.6%) provide the complete and necessary information required for reproducibility such as prompts and environment.
4.4 Guideline for Evaluation Analysis
Figure 6: Guideline for Evaluation Analysis
Explanation – We listed 10 criteria for the evaluation analysis phase, as shown in Figure 6. Regarding the perspectives of analysis, inspired by classic measurement theory Suppes et al. (1962), we suggest four essential perspectives, including difficulty (whether a benchmark is appropriately challenging for LLMs), stability (whether the results are consistent through repeated trials), differentiability (whether benchmarks can differentiate the strengths and weaknesses of various LLMs).
Moreover, effective presentation of results using clear visual and textual descriptions is recommended to ensure the findings are understandable and actionable. The phase concludes with the suggestion to interpret and explain the results comprehensively, providing a basis for future research and application enhancements.
Key Statistics – Because experimental analysis is relatively subjective and cannot be obtained through mechanical scanning, we focus on 30 representative focus benchmarks (Appendix C), covering the highest cited and latest benchmarks in the five tasks (Section 3.2 Step 3). Figure 37 shows an example from CruxEval Gu et al. (2024) where the experimental scores can hardly be read from the figures.
Also, explaining experiment results is crucial for other practitioners to understand what the outcomes mean in the context of the research questions. According to our statistics (Appendix A.5), 70% benchmarks have detailed explanations and analyses of their evaluation results, while still 30% have not. Indeed, an explanation contributes to the body of knowledge by making it possible to understand and compare results with previous studies, promoting transparency within the community.
{mdframed}
[style=MyFrame] ▲ Severity – The analysis of experimental data and the clarity of data presentation may receive less attention and worth consideration.
4.5 Guideline for Benchmark Release
Figure 7: Guideline for Benchmark Release
Explanation – Finally, releasing a benchmark for open access also needs careful consideration. We offered 10 suggestions for this step, as shown in Figure 7, to highlight essential steps for public release preparation, emphasizing accessibility and ethical compliance. This includes setting an appropriate license to clarify usage rights, conducting a thorough review to eliminate sensitive or harmful content such as the API keys to access LLMs, the personal emails or toxic code comments Miller et al. (2022) unless they are a part of the benchmark, and ensuring transparency and reproducibility by making all related materials openly available. Detailed prompts and clear descriptions of the experimental setup are advised to facilitate replication. Additionally, providing user manuals and evaluation interfaces is crucial for effective user engagement with the benchmark, enhancing its reliability and value for the research community.
Key Statistics – The final step involves the release of the benchmark. The fundamental requirement for releasing a benchmark is that it must be open-sourced. However, surprisingly, we observed that 5.1% of the benchmarks are only partially open-sourced (e.g., missing some subjects or tests), and 5.8% are not open-sourced at all (e.g., links/web pages are no longer active). 19.3% have not properly set up the license. Furthermore, prompts, which are necessary for reproducibility, are not disclosed in 52.6% of the benchmarks (Figure 39). Not to mention the lack of public information on experimental settings and experimental parameters (Figure 43). Furthermore, only 16.7% of the benchmarks make their logged experimental results publicly available. See Appendix A.6 for details.
{mdframed}
[style=MyFrame] ▲ Severity – The release of existing benchmarks exhibits several issues. For example, over 10% of the benchmarks are either not open to public access or are only partially open-sourced. Only 47.4% of benchmarks are released with replicable prompts.
5 Human Study
To delve deeper into the integration of knowledge and action, we surveyed 49 global researchers in AI (42.6%) and SE (57.14%), as shown in Figure 50. Each participant had published at least one research paper, and about half had constructed code-related benchmarks. See Appendix B.
First, all participants agreed that having a checklist for benchmark construction would contribute to the quality of the benchmark. 47/55 criteria in How2Bench are deemed important by more 80% participants. Additionally, among the 21 participants who have constructed code-related benchmarks, 53 out of 55 criteria were deemed important by all benchmark developers; only two criteria (criteria 3 and 4 in Section 4) were considered unimportant by a few individuals (3 and 2 participants, respectively). Additionally, we received two valuable suggestions that draw importance to recording the time/monetary costs of constructing the benchmark and conducting the experiments.
However, we also identified some notable gaps in awareness. First, regarding the data preparation, more than 15% of participants were not aware that the selection of data should consider the target scope of the evaluation set (i.e., the data must be representative), and 16% of participants were unaware of the need for data denoising. This oversight can significantly affect the validity and generalizability of experimental results, underscoring the importance of a comprehensive understanding of data handling for reliable research outcomes. Second, regarding evaluation replicability and reliability. Over 40% of participants believe that recording and publicizing the hardware and software environments, software versions, and libraries used in experiments is not important, with more than 20% still considering it unimportant despite already done so. This reveals a significant lack of awareness about the impact that experimental environments can have on the reliability, reproducibility, and stability of evaluation results. In fact, various studies have demonstrated that different experimental environments, parameters, and prompts can lead to substantial variations in outcomes Xiao et al. (2024); Wang et al. (2019, 2023a).
6 Conclusion
This paper proposes a rigorous guideline consisting of 55 checklists covering the benchmark development lifecycle. After investigating over 270 code-related benchmarks, we exposed their merits and limitations and provided suggestions for improving them. Finally, our human study reveals the neglect of details that may affect the benchmark’s reliability. In the long run, How2Bench helps to improve the overall quality of benchmarks in the community due to the propagation among benchmarks.
Limitations
This paper has two primary limitations that offer avenues for future research. First, the collection of code-related benchmarks may be incomplete. To minimize this limitation, we covered papers published over the last decade, and conducted multiple rounds of snowballing. Ultimately, we collected 274 benchmarks, which is comparable to the number included in recent surveys Hou et al. (2023); Schäfer et al. (2024) in the field. Second, the study involved substantial manual analysis, which could lead to oversight and discrepancies in the statistical results. To mitigate this issue, we ensured that each benchmark was double-checked by at least two authors and underwent multiple rounds of iteration. Third, the guidelines may not cover all the details. Constructing a code-related benchmark involves numerous details, and some criteria are task-specific. To overcome this limitation, we iteratively refined the guidelines, interviewed practitioners, and tried to cover the entire benchmark development process as thoroughly as possible. Last, the human study participants may exhibit subjectivity. To address this limitation, we endeavored to include a broad range of practitioners and seasoned researchers with experience in both AI and SE, aiming for the least biased results possible.
Appendix A Statistics of studied benchmarks
In this section, we conducted a comprehensive and detailed statistical analysis of the 274 benchmarks collected.
Figure 8: Relationships between Benchmarks
A.1 Profile of Studied Benchmarks
We first show the trend in the development of benchmarks from 2014 to 2024. As shown in Figure 9, the data shows a modest beginning, with only a handful of benchmarks created annually until 2017. From 2018 onwards, there is a noticeable uptrend in benchmark creation, culminating in a significant jump to 149 benchmarks in 2024. This sharp increase indicates a recent heightened interest and demand for comprehensive code-related benchmarks for LLMs, reflecting the evolving complexities and expanding requirements of automated software engineering.
Hierarchy of Benchmarks. Figure 8 visualize the inheritance relationships among benchmarks, indicating that the benchmarks on the left serve as sources for those on the right. It highlights that 18% (50 out of 274) of benchmarks act as data sources, continuously benefiting the construction of subsequent benchmarks.
Figure 8 reveals that HumanEval Chen et al. (2021a), as the most significant source benchmark, benefits at least 15 downstream benchmarks, followed by MBPP Austin et al. (2021) and CodeSearchNet Husain et al. (2019). From the right side of the figure, some benchmarks, like VulBench Gao et al. (2023b), incorporate methodologies or data from 4 previous benchmarks, and codeRagBench Wang et al. (2024d) integrates elements from 8 prior benchmarks.
This hierarchical structure among benchmarks also alerts us that the data quality of a benchmark not only affects its own credibility but can continue to impact others if it serves as a source. This underscores the importance of adhering to stringent guidelines during benchmark development and highlights the crucial role of establishing standards to ensure the integrity and utility of benchmark data across research and development efforts.
Figure 9: Benchmark Distribution over Years
Annual Trend. Regarding the coding tasks, Figure 11 illustrates the distribution of various coding tasks across benchmarks. It is clear that the task of Code Generation is most prevalent, with 99 benchmarks focusing on this area according to 36% (99/274) of studied benchmarks, indicating a significant interest in generating code automatically. Program Repair and Defeat Detection are well-represented, with 27 and 25 benchmarks, respectively, reflecting the importance of correcting code and detecting defects.
Citation distribution. We also visualized the citations of 274 code-related benchmarks. The citation statistics were collected on September 1st, 2024. From Figure 10, we can see a clear long-tail trend of the citations, from the highest 2735 (HumanEval Chen et al. (2021a)) to the lowest 0.
Figure 10: Citation Distribution of Benchmarks
Coding Task. Tasks like Code Summarization and Text2SQL are similarly significant, each with 25 and 22 benchmarks. These tasks focus on making code more understandable and converting natural language queries into SQL queries. Other tasks, such as Code Retrieval, Code Reasoning, and Code Translation, are represented with 18, 17, and 16 benchmarks, respectively. Lesser-represented benchmarks are Test Generation, Code Optimization, and Code Completion, each represented by 8 and 7 benchmarks, indicating the inadequacy of these tasks.
Figure 11: Benchmark Distribution over Tasks
Programming Languages. Figure 12 shows the distribution of benchmarks across various programming languages. The overall trend indicates a strong preference for benchmarking Python, which leads with 158 benchmarks, followed by Java and C++, with 107 and 63, respectively. The graph also reveals a diverse range of languages being used. In total, 724 programming languages are studied by these 274 benchmarks. Though some programming languages, such as Kotlin, Swift, and Scala, are less frequently exercised, the benchmarks involving them are tailored to different application needs and technology environments. This distribution shows the existing benchmarks are dominated by three mainstream programming languages, leaving other programming languages less studied and benchmarked.
Figure 12: Benchmark Distribution over Programming Language
Natural Language. Figure 13 illustrates the distribution of benchmarks for different natural languages. The bar chart overwhelmingly shows that English is the dominant language, with 192 benchmarks highlighting its ubiquity in research and development. Other languages have significantly fewer benchmarks, with six for Chinese and only two each for Japanese, Russian, and Spanish. The category labeled “Other” includes 20 benchmarks spread across other natural languages, indicating some diversity but limited attention to non-English benchmarks. This distribution highlights the prominence of English in the global research community and also demonstrates the uneven representation of natural languages in the studied benchmarks.
Figure 13: Benchmark Distribution over Natural Language
Modals in the benchmarks. Figure 14 presents the distribution of benchmarks according to the type of language used in their prompts. The chart shows that the majority, at 47.1%, of the benchmarks use a combination of natural language and programming Language, followed by PL only (31.0%) and NL only (21.9%).
Figure 14: Benchmark Distribution over Modal in Prompt
Granularity. The code snippet in a code-related benchmark varies from statement-level (i.e., one line of code. For example, CoNaLa Yin et al. (2018) and Math-QA Amini et al. (2019)), function-level (i.e., a function unit of code. For example, HumanEval Chen et al. (2021a) and MBPP Austin et al. (2021)), class-level (i.e., a class with multiple function units of code. For example, ClassEval Du et al. (2023b)) and project-level (i.e., a project with multiple classes or modules. For example, DevEval Li et al. (2024a) and JavaBench Cao et al. (2024a)). Figure 15 illustrates the granularity levels at which benchmarks are typically conducted. The chart shows that the majority of benchmarks, comprising 71.8%, focus on the function level. Projects constitute 15.0% of the benchmarks. Class-level granularity is the least represented at 2.6%.
{mdframed}
[style=MyFrame] The majority of benchmarks are currently at the function level (70+%), followed by the project level (15+%). This indicates that the current major demand is for assessing individual functions within a single task, followed by the demand for evaluating functionalities more aligned with actual project-level code development.
Figure 15: Benchmark Distribution over Granularity
A.2 Statistics about Benchmark Design
Design of Studied Capabilities. To understand whether benchmark developers recognize the capabilities of LLMs they aim to evaluate, we carefully analyzed 30 representative benchmarks (Appendix C) to see if they clearly specify the capabilities being assessed by their benchmarks. As shown in Figure 16, 90% of benchmarks explicitly specify the capabilities (e.g., intention understanding, problem solving, testing, debugging capabilities)to be evaluated, while 10% do not. The statistics show that the most highly cited benchmarks clearly define the assessment capabilities.
Figure 16: Benchmark Distribution Over Capabilities Consideration
Furthermore, we investigated the 30 focused benchmarks and identified a case (Figure 17) from MBPP Austin et al. (2021) where the case is likely to fall outside of the targeted capability of the benchmark. In particular, MBPP Austin et al. (2021) aims to “measure the ability of these models to synthesize short Python programs from natural language descriptions” for “entry-level programmers”. As we can see from Figure 17, the prompt requires LLMs to “Write a function to calculate the dogs’ years.” Simply from this description, an entry-level programmer is unlikely to write a correct program without knowing the conversion equation from dogs’ year to dogs’ age. In other words, this case is more about assessing whether LLMs have acquired this specific knowledge rather than evaluating the most fundamental programming skills.
Figure 17: An Example of Out-of-capability Case from MBPP Austin et al. (2021).
Design of Studied Application Scenarios. Similarly, to understand whether benchmark developers scoped the application scenarios of LLMs they aim to evaluate, we carefully analyzed 30 representative benchmarks (Appendix C) to see whether they explicitly specify the application scenarios their benchmarks target. As shown in Figure 18, 70% representative benchmarks have clearly specified application scenarios (e.g., programming assistant), while the rest do not. Indeed, clearly defining the application scenarios could help benchmark constructors establish precise goals for the design and development of the benchmark, ensuring accuracy in the evaluation.
Figure 18: Benchmark Distribution Over Expected Application Scenario Consideration
A.3 Statistics about Data Preparation
A.3.1 Data Preprocessing
Data Deduplication. During benchmark preparation, data cleaning and preprocessing are necessary. However, as shown in Figure 19, only 38% benchmarks have deduplicated the collected data. More than half of them didn’t mention this process.
Figure 19: Benchmark Distribution over Deduplication
To investigate the situation, we went through the 30 representative benchmarks (Listed in Appendix C) and found two duplicated subjects in MBPP Austin et al. (2021). Tasks with id 71 and 141 examined the same functionality, i.e., “Write a function to sort a list of elements.”, collected from the same source.
Figure 20: A Counterexample of Rule 16 from MBPP Austin et al. (2021).
{mdframed}
[style=MyFrame] The significance of data preprocessing, such as deduplication, is frequently overlooked by benchmark builders, leading to data duplication even in highly cited benchmarks.
Data Quality Assurance. Ensuring data quality for the benchmark is essential. However, our statistics (Figure 21) show disappointing results. 67.9% of benchmarks do NOT take any measures for data quality assurance. Among those benchmarks that do incorporate data quality measures, the majority rely on manual checks, which accounts for 22.6%. Other countermeasurements, such as code execution, constitute only 2.2%, while verification using LLMs accounts for 1.5%. Additional methods, such as using download counts as a basis, are also employed.
Figure 21: Benchmark Distribution over Quality Assurance Method
Additionally, we dived into the 30 representative benchmarks (Listed in Appendix C) and identified an example where the code cannot be executed successfully. As shown in Figure 22, the function swap() in line 7 has not been defined, so the execution of the code would fail if the code has been executed. This highlights a significant gap in ensuring the reliability and validity of benchmark data, underscoring the need for more rigorous and automated data quality assurance practices.
Figure 22: An Example from MBPP  Austin et al. (2021) that failed to be executed.
Data Contamination Resolution. Data contamination Golchin and Surdeanu (2023); Cao et al. (2024b) threat has been widely discussed. A benchmark with contaminated data may yield overclaimed results, misleading the understanding of the LLMs’ capabilities. According to our statistics (Figure 23 on benchmarks from the year 2023 to 2024 (the duration when most LLMs were launched), most (78 %) benchmarks were not aware of and have not taken any measures to alleviate data contamination, being vulnerable to data contamination threat.
Figure 23: Benchmark Distribution over Quality Assurance on Data Contamination
A.3.2 Statistics about Data Curation
Ground truth solutions. Figure 24 shows that although the majority (92.3%) of benchmarks provide reference code as ground truth, there are 5% of benchmarks without reference code. Although it is not compulsory as long as object measurements (e.g., test cases) are provided, a reference code is still recommended. Indeed, if a benchmark provides reference code, its reliability tends to be better because it ensures that there are feasible solutions for the tasks involved. This guarantees that the tasks are theoretical and practically solvable, enhancing the benchmark’s usefulness and credibility in real-world applications.
Figure 24: Benchmark Distribution over Solution
Additionally, the correctness of the ground truth solution should also be noted. Figure 25 shows an incorrect code solution provided in HumanEval Chen et al. (2021a). This should draw benchmark constructors’ attention to the correctness of the benchmark reference code.
Figure 25: An Example from Humaneval  Chen et al. (2021a) which shows an incorrect solution provided in the benchmark.
Oracle. An oracle Barr et al. (2014) is a way to determine whether the output is correct or not. For example, assume the output of LLMs is in the form of code, then an oracle could be running tests against the code and see whether the code can pass all the tests. Figure 26 shows the distribution of types of oracle that are used in these benchmarks. Clear that the exact match 41.97% (115/274) and test case passing (114/274) 41.6% are the most common oracle used in code-related benchmarks, followed by thresholds (i.e., similarities smaller than a specific threshold).
Figure 26: Benchmark Distribution over Test Orcale
Code test coverage Ivanković et al. (2019), as a common oracle for code-related benchmarks, has been widely adopted to determine the output correctness. It should be considered if a benchmark uses test case passing as a criterion for the correctness of the generated code. Otherwise, a test could be too weak to detect the existence of a defect in the generated code. For example, as pointed out by prior work Liu et al. (2023a), existing benchmarks such as HumanEval Chen et al. (2021a) and MBPP Austin et al. (2021) still suffer from “insufficient tests”, allowing incorrect code to pass all the tests without capturing the bugs.
Despite its importance, as shown in Figure 27, among the benchmarks that use test cases as the oracle, only 8.7% considered and reported “test coverage” explicitly in their papers, while 87.8% did not mention the test coverage of their provided code.
Figure 27: Benchmark Distribution over Test Coverage
Furthermore, we dived into 30 representative benchmarks (Listed in Appendix C) and identified an example (Figure 28) from MBPP Austin et al. (2021) where the test is incorrect. It alerts us that both the quality of the test and the test adequacy (e.g., code coverage) should be considered.
Figure 28: An Example of Incorrect Tests from MBPP Austin et al. (2021).
A.4 Statistics about Evaluation
Studied LLMs. We summarize the number of LLMs that have been evaluated in each benchmark evaluation. As shown in Figure 29, the majority of benchmarks are tested with a smaller number of LLMs, as evidenced by the bars for 1 to 9 LLMs. Interestingly, there is a significant peak at the 10-49 LLM range, with 56 benchmarks falling into this category, suggesting a common practice of testing benchmarks extensively but not excessively. The drop in benchmarks tested with 50-99 and 100+ LLMs indicates that very extensive testing across numerous models is less common, potentially due to resource constraints or diminishing returns on insights gained from additional models.
Figure 29: Benchmark Distribution over LLM Experimented
Additionally, we listed the top-10 LLMs by the number of code-related benchmarks they have been evaluated, as shown in Figure 30. GPT series leads significantly with 116 benchmarks, suggesting its widespread adoption and possibly its versatility or superior performance in handling code-related tasks. The rest, including CodeLlama, StarCoder, CodeGen, and others, show varying degrees of involvement, with numbers ranging from 60 down to 24 benchmarks for Claude. This figure may provide a reference for choosing which model to evaluate. In addition, it is worth mentioning that different LLMs should be considered considering different coding tasks.
Figure 30: Top-10 Studied LLMs for Code-related Benchmarks
Experiment Environments. The experimental environment (such as the operating system and hardware) is important for the reproduction of the experiment. However, Figure 32 and Figure 31 highlight a significant gap. A mere 27.4% of benchmarks document the devices used in their experiments, leaving a substantial 72.6% that do not. The situation appears even more dire when considering os, with only 3.6% of benchmarks documenting the OS used, while a staggering 96.4% neglect to record this information.
Figure 31: Benchmark Distribution over Recording Experiment Devices
Figure 32: Benchmark Distribution over Recording Experiment OS
Prompting and Prompting Strategies Prompting has a direct impact on the quality of the LLMs’ output results Wei et al. (2022); He et al. (2024a); Jin et al. (2024). So, we summarized whether different prompting strategies have been evaluated and statistics the distribution. Figure 33 shows the usage of four kinds of prompts: zero-shot, few-shot, chain-of-thought, and retrievals (RAG). From Figure 33, we can see that a vast majority (94.9%) benchmarks were evaluated in a zero-shot context setting, while only 21.2% benchmarks were evaluated in a few-shot manner. Even fewer benchmarks were evaluated under the COT and RAG settings, utilized by only 8.8% and 2.6%.
Figure 33: Benchmark Distribution over Context Setting
Prompt Quality The prompt quality also greatly impacts the LLM evaluation He et al. (2024b). So, carefully designing a prompt needs consideration. However, as shown in Figure 34, 73.3% representative benchmarks (Appendix C) do not validate whether the prompt they used is well-designed.
Figure 34: Benchmark Distribution Over Validation of Prompts
Repeated Experiment Given the random nature of LLMs, the experiments are expected to repeat, ensuring the stability and reliability of the results. However, as shown in Figure 35, only 35.4% benchmarks went through a repeated experiment, while a majority of 64.6% opted against repeating the experiment. This reflects a lack of awareness regarding the stability and reproducibility of evaluations.
Figure 35: Benchmark Distribution over Repeating the Experiment
A.5 Statistics about Analysis
Experiment Explanation. Explaining experiment results is crucial for other practitioners to understand what the outcomes mean in the context of the research questions. So, we investigate whether the representative benchmarks (Appendix C) have explained the experiment results. As shown in Figure 36, 70% benchmarks have detailed explanations and analyses of their evaluation results, while still 30% have not. Indeed, an explanation contributes to the body of knowledge by making it possible to understand and compare results with previous studies, promoting transparency within the community.
Figure 36: Benchmark Distribution Over Explaining the Experiment
A clear and precise presentation of experimental results is important for enabling robust interpretation and comparison across benchmarks. However, further examination of the 30 representative benchmarks (listed in Appendix C) revealed notable deficiencies in result visualization. As shown in Figure 37, CruxEval Gu et al. (2024) exhibits unclear experimental result presentation. Specifically, the scatter plot suffers from ambiguous labeling, poor readability of axis values, and inconsistent marker encoding, making it difficult for researchers to extract meaningful insights. Such presentation shortcomings obscure the performance relationships between methods and compromise the benchmark’s usability for fair evaluation. To address these issues, benchmarks should adopt standardized and well-documented visualization practices, ensuring results are interpretable, accessible, and reproducible.
Figure 37: An Example of Unclear Experiment Analysis and Display from CruxEval Gu et al. (2024)
A.6 Statistics about Release
Data Accessibility. The fundamental requirement for releasing a benchmark is that it must be open-sourced. However, surprisingly, as shown in Figure 38, we observed that 5.1% of the benchmarks are only partially open-sourced (e.g., missing some subjects or tests), and 5.8% are not open-sourced at all (e.g., links/web pages are no longer active).
Figure 38: Benchmark Data Availability
Prompt Accessibility. Detailed prompts are essential for ensuring the reproducibility and transparency of code-related benchmarks. However, as shown in Figure 39, we found that 52.6% of benchmarks do not provide detailed prompts, limiting the ability to accurately replicate and evaluate the performance of LLMs. This lack of prompt disclosure highlights a gap in benchmark design practices, as prompts are often indispensable for understanding model performance under specific conditions. While 47.4% of benchmarks include such prompts, the absence of comprehensive prompt documentation in over half of the cases raises concerns about the consistency and reproducibility of reported results.
Figure 39: Availability of Prompts
Logging Info Accessibility. Providing detailed logging information, including comprehensive experimental results, is essential for ensuring transparency, verifiability, and reproducibility in benchmarking research. However, as shown in Figure 40, only 16.7% of the benchmarks make their experimental results publicly available, while 80.0% fail to disclose this critical information. Alarmingly, an additional 3.3% provide only partial logging details, further complicating result verification. The absence of complete logging information creates significant barriers for researchers attempting to reproduce experiments or validate reported findings, thereby undermining the reliability of benchmarks. To address this, we emphasize the necessity of making detailed logging information, including intermediate results and metrics, publicly accessible to uphold rigorous scientific standards and foster trustworthy comparisons across models.
Figure 40: Availability of Logging Information
User Manual Accessibility. A high-quality user manual, such as a well-documented README file, is crucial for enhancing benchmark usability, enabling users to understand the dataset, execute provided scripts, and reproduce results efficiently. However, our analysis revealed that a significant number of benchmarks lack comprehensive user manuals, hindering accessibility and adoption. As depicted in Figure 41, poorly structured or incomplete manuals often omit essential components such as benchmark introductions, usage instructions, and evaluation scripts. This creates unnecessary barriers for researchers who rely on these manuals for setup and experimentation. To address this, we advocate for benchmarks to include clear, standardized user manuals that provide an overview of the benchmark, step-by-step execution guides, and troubleshooting instructions, ensuring a seamless and reproducible user experience.
Figure 41: Availability of User Manual
Convenient Evaluation Interface Availability.
Providing convenient evaluation interfaces is essential for enhancing the usability and accessibility of benchmarks, enabling researchers to easily reproduce results and compare models. As shown in Figure 42, 16.7% of benchmarks fail to offer any evaluation interfaces, imposing significant barriers to usability. While a majority of benchmarks (83.3%) provide such interfaces, including command-line tools, Docker images, or scripts, the absence of standardized and user-friendly evaluation tools in a notable minority of cases highlights an area for improvement. Benchmarks without convenient evaluation interfaces require users to spend additional effort in setup and result verification, which can discourage adoption and hinder reproducibility. To address this, we emphasize the importance of releasing benchmarks with well-documented, ready-to-use evaluation pipelines to promote efficient, reliable, and fair benchmarking practices.
Figure 42: Availability of Convenient Evaluation Interfaces
Temperature Records. One critical parameter for benchmarking is the temperature setting, which influences stochasticity in LLMs. As shown in Figure 43, we observed that 57.3% of benchmarks fail to record the temperature setting, hindering reproducibility and fair evaluation. While 42.7% of benchmarks do document this parameter, the majority omission highlights an overlooked yet essential aspect of benchmark transparency.
Figure 43: Benchmark Distribution over Recording Temperature
License Provision. Releasing benchmarks under a clear and accessible license is fundamental for legal compliance and ensuring open collaboration. Figure 44 reveals that 19.3% of benchmarks do not provide a license, limiting their usability and distribution. Encouragingly, 80.7% of benchmarks do include a license, but the lack of licensing in nearly one-fifth of the benchmarks raises concerns about widespread adoption and usage. These findings emphasize the need for standardized practices in benchmark releases to promote legal clarity and accessibility.
Figure 44: Provision of License
Data Security. Ensuring data security is a critical yet often overlooked aspect of benchmark development. Sensitive information, such as API keys, credentials, or private tokens, should never be included in benchmark releases. However, further investigation into 30 representative benchmarks (listed in Appendix C) revealed instances of sensitive data leakage. As shown in Figure 45, XSemPLR Zhang et al. (2023b) inadvertently included an API key in its release, a critical oversight that can expose resources to external exploitation.
Figure 45: An Example of API Key Leakage in Benchmark Release from XSemPLR Zhang et al. (2023b).
Similarly, Figure 46 highlights an example from CrossVul Nikitopoulos et al. (2021), where personal names and email addresses were unintentionally disclosed. Such leakage poses risks of unauthorized access and resource misuse, potentially compromising systems and research integrity.
Figure 46: An Example of Name & Email Leakage in Benchmark Release from CrossVul Nikitopoulos et al. (2021).
Usability. Clear and comprehensive documentation is crucial for ensuring the usability of benchmarks, as poorly written instructions can significantly hinder adoption and reproducibility. We dived into the 30 representative benchmarks (listed in Appendix C) and identified an example where the README file provided insufficient and unclear information. As shown in Figure 47, VulDeePecker Li et al. (2018b) includes a less-than-ideal ReadMe file, which lacks essential usage instructions and evaluation guidelines, making the benchmark difficult to understand and deploy. In contrast, Figure 48 highlights APPS Hendrycks et al. (2021), which provides well-structured and easy-to-follow documentation. The APPS benchmark includes step-by-step instructions for generating, evaluating, and analyzing results, enabling users to efficiently reproduce experiments. These observations emphasize the importance of high-quality documentation for benchmarks to enhance accessibility, reduce friction in usage, and foster reproducible research.
Figure 47: An Example of Unreadable and Hard-to-Use README in Benchmark Release from VulDeePecker Li et al. (2018b).
Figure 48: A Good Example of Easy-to-Read README in Benchmark Release from APPS Hendrycks et al. (2021).
Appendix B Details of Human Study
B.1 Interviewee Selection
The selection of interviewees is pivotal to ensuring the representativeness and relevance of the data collected. This involves identifying individuals with the knowledge or experience pertinent to the research theme.
To this end, we chose graduate students from SE or AI fields who have published at least one paper. This criterion ensures that participants have research experience and judgment capabilities. The focus on SE and AI fields is due to their likely interest in code benchmarks. Particularly, we aimed to recruit individuals who have published papers on code benchmarks to obtain firsthand feedback from experienced benchmark developers.
B.2 Survey Question Design
Questions. The body of the survey was divided into five stages of benchmark development (following Figure 1), with necessary background information provided for each stage. Each criterion in How2Bench was slightly modified to be in the first-person perspective, making it easier for interviewees to empathize and answer the questions from their own viewpoint. Finally, to facilitate comprehension, questions and instructions were translated into both English and Chinese.
Question Setting. To minimize the effort required from respondents, we designed single-choice questions with four options:
❒ I found it important, and I have done it.
❒ I found it important, although I haven’t done it.
❒ I found it not important, but I have done it.
❒ I found it not important, and I wouldn’t do it.
This format is intended to orthogonally explore the correlation between awareness and behavior.
B.3 Interview Process
Questionnaire Distribution The questionnaire was distributed via online platforms, targeting academic and professional networks related to SE and AI. The distribution started on October 27, 2024, and ended on November 27th, 2024, lasting one month.
Results Collection The responses were automatically collected through the online platform used for distribution.
Survey Screening Since the requirement was for participants who have published papers, responses from those selecting "No" to having published a paper were excluded. Additionally, incomplete surveys where not all questions were answered were also considered invalid and excluded from the analysis.
B.4 Interview Result Analysis
In total, we collected 50 responses. The respondents were from seven regions, including the United States, the United Kingdom, Germany, Australia, China, and others, as shown in Figure 49. Only one survey was invalid due to the respondent selecting “have not published a paper”, leaving 49 valid surveys for analysis. A breakdown of the respondents’ demographics is shown in Figure 50. And the detailed response for all 55 criteria in How2Bench are shown in Figure 51 and Figure 52.
Figure 49: Geographical Distribution of Interviewees
Figure 50: Demography of Interviewees
Figure 51: Results of Human Study (Questions 1 - 28
Figure 52: Results of Human Study (Questions 29 - 55
Appendix C List of Studied Benchmarks (Focused Ones)
Code Generation: Five with top citations:
  • HumanEval Chen et al. (2021a)
  • MBPP Austin et al. (2021)
  • CodeContest Li et al. (2022)
  • leetcodehardgym Shinn et al. (2023)
  • APPS Hendrycks et al. (2021)
The latest one as of 31/8/2024:
  • VerilogEval Pinckney et al. (2024)
Defect Detection: Five with top citations:
  • VulDeePecker Li et al. (2018b)
  • Devign Zhou et al. (2019)
  • Chromium and Debian Chakraborty et al. (2022)
  • $\mu$VulDeePecker Zou et al. (2020)
  • Synthetic Dataset Hellendoorn et al. (2020)
The latest one as of 31/8/2024:
  • VulDetectBench Liu et al. (2024c)
Program Repair: Five with top citations:
  • Defects4J Just et al. (2014)
  • BFP Tufano et al. (2019)
  • MANYBUGS, INTROCLASS Le Goues et al. (2015)
  • HumanEval-Java Jiang et al. (2023)
  • QuixBugs Prenner et al. (2022)
The latest one as of 31/8/2024:
  • DebugBench Tian et al. (2024)
Code Summarization: Five with top citations:
  • CODE-NN Iyer et al. (2016)
  • Java-small/med/large Alon et al. (2019)
  • code-summarization-public Wan et al. (2018)
  • HumanEvalPack Muennighoff et al. (2024)
  • Shrivastava et al. Shrivastava et al. (2023b)
The latest one as of 31/8/2024:
  • Long Code Arena Bogomolov et al. (2024)
Text To SQL: Five with top citations:
  • WikiSQL Zhong et al. (2017)
  • Spider Yu et al. (2018)
  • Advising Finegan-Dollak et al. (2018)
  • BIRD Li et al. (2023a)
  • Spider-Realistic Deng et al. (2021)
The latest one as of 31/8/2024:
  • AMBROSIA Saparina and Lapata (2024)
Appendix D List of Studied Benchmarks (Full)
We collected and studied 274 code-related benchmarks. We then listed and grouped them by year.
2024:
  • CodeEditorBench Guo et al. (2024)
  • MHPP Dai et al. (2024)
  • LiveCodeBench Jain et al. (2024)
  • CodeAgentBench Zhang et al. (2024a)
  • CruxEval Gu et al. (2024)
  • BigCodeBench Zhuo et al. (2024)
  • OOPEval Wang et al. (2024b)
  • DevEval Li et al. (2024a)
  • Long Code Arena Bogomolov et al. (2024)
  • CodeRAGBench Wang et al. (2024d)
  • ScenEval Paul et al. (2024)
  • AICoderEval Xia et al. (2024b)
  • VersiCode Wu et al. (2024b)
  • VHDL-Eval Vijayaraghavan et al. (2024)
  • NaturalCodeBench Zhang et al. (2024b)
  • CodeGuard+ Fu et al. (2024)
  • PECC Haller et al. (2024)
  • USACO Shi et al. (2024b)
  • ParEval Nichols et al. (2024)
  • MxEval Athiwaratkun et al. (2022)
  • MMCode Li et al. (2024c)
  • Plot2Code Wu et al. (2024a)
  • ChartMimic Shi et al. (2024a)
  • DebugBench Tian et al. (2024)
  • PythonIO Zhang et al. (2024c)
  • StaCCQA Yang et al. (2024a)
  • RepoQA Liu et al. (2024a)
  • PRIMEVUL Ding et al. (2024a)
  • VulDetectBench Liu et al. (2024c)
  • ProCQA Li et al. (2024e)
  • CoSQA+ Gong et al. (2024)
  • JavaBench Cao et al. (2024a)
  • HumanEvo Zheng et al. (2024)
  • REPOEXEC Hai et al. (2024)
  • EHR-SeqSQL Ryu et al. (2024)
  • BookSQL Kumar et al. (2024)
  • AMBROSIA Saparina and Lapata (2024)
  • WUB, WCGB Yun et al. (2024)
  • RES-Q LaBash et al. (2024)
  • PythonSaga Yadav et al. (2024b)
  • Mercury Du et al. (2024)
  • ENAMEL Qiu et al. (2024b)
  • RealHumanEval Mozannar et al. (2024)
  • CoderUJB Zeng et al. (2024)
  • EvoEval Xia et al. (2024a)
  • ML-Bench Liu et al. (2023c)
  • VerilogEval Pinckney et al. (2024)
  • CodeApex Fu et al. (2023)
  • HumanEvalPack Muennighoff et al. (2024)
  • HumanEval+ Liu et al. (2023b)
  • HumanEval-X Zheng et al. (2023a)
  • XCodeEval Khan et al. (2024)
  • CoderEval Yu et al. (2023)
  • CodeXGLUE Lu et al. (2021)
  • VulnPatchPairs Risse and Böhme (2024)
  • WikiSQL Zhong et al. (2017)
  • CrossCodeEval Ding et al. (2023)
  • SWE-bench Jimenez et al. (2024)
  • BAIRI et al. Bairi et al. (2024)
  • BioCoder Tang et al. (2024)
  • RepoBench Liu et al. (2024b)
  • NoFunEval Singhal et al. (2024)
  • CoCoMIC Ding et al. (2024b)
  • Java-small/med/large Alon et al. (2019)
  • FixEval Haque et al. (2023)
  • CommitBench Schall et al. (2024)
  • InfiAgent-DABench Hu et al. (2024)
  • InfiBench Li et al. (2024d)
  • Design2Code Si et al. (2024)
  • MatPlotBench Yang et al. (2024b)
  • EditEval Li et al. (2024b)
  • D1, D2, D3 Huang et al. (2024)
  • RepoEval Liao et al. (2024)
  • BetterTypes4Py, InferTypes4Py Wei et al. (2023)
  • HumanEval-Java Jiang et al. (2023)
  • PIE Shypula et al. (2024)
  • EvalGPTFix Zhang et al. (2023a)
  • EHRSQL Lee et al. (2023)
  • Spider2-V Cao et al. (2024c)
  • TESTEVAL Wang et al. (2024c)
  • ChatTester Yuan et al. (2023b)
  • Code Lingua Pan et al. (2024)
  • EffiBench HUANG et al. (2024)
  • CRUXEval-X Xu et al. (2024)
  • DomainEval Zhu et al. (2024)
2023:
  • MCoNaLa Wang et al. (2023b)
  • MultiPL-E Cassano et al. (2022)
  • ODEX Wang et al. (2022)
  • TACO Li et al. (2023b)
  • DOTPROMPTS, MGDMICROBENCH Agrawal et al. (2023)
  • StudentEval Babe et al. (2024)
  • CodeTransOcean Yan et al. (2023)
  • G-TransEval Jiao et al. (2023)
  • AVATAR Ahmad et al. (2023)
  • RunBugRun Prenner and Robbes (2023)
  • VulBench Gao et al. (2023b)
  • DiverseVul Chen et al. (2023)
  • Hellendoorn et al. Hellendoorn et al. (2020)
  • XSemPLR Zhang et al. (2023b)
  • BIRD Li et al. (2023a)
  • Stack-Repo Shrivastava et al. (2023a)
  • RepoEval Liao et al. (2024)
  • MTPB Nijkamp et al. (2022)
  • ARCADE Yin et al. (2023)
  • Shrivastava et al. Shrivastava et al. (2023b)
  • Grag et al. Garg et al. (2022)
  • GSM-HARD Gao et al. (2023a)
  • InferredBugs Jin et al. (2023)
  • LeetcodeHardGym Shinn et al. (2023)
  • APIBench Patil et al. (2023)
  • ClassEval Du et al. (2023b)
  • CommitChronicle Eliseeva et al. (2023)
  • TeCo Nie et al. (2023)
  • TESTPILOT Schäfer et al. (2024)
2022:
  • AixBench Hao et al. (2022)
  • TypeBugs Oh and Oh (2022)
  • XLCoST Zhu et al. (2022)
  • CS1QA Lee et al. (2022)
  • Chromium and Debian Chakraborty et al. (2022)
  • Spider-Realistic Deng et al. (2021)
  • Spider-SS Gan et al. (2022)
  • DSP Chandel et al. (2022)
  • CodeContest Li et al. (2022)
  • PandasEval, NumpyEval Zan et al. (2022b)
  • TorchDataEval, MonkeyEval, BeatNumEval Zan et al. (2022a)
  • DS-1000 Lai et al. (2023)
  • MCMD Tao et al. (2022)
  • ExeDS Huang et al. (2022)
  • QuixBugs Prenner et al. (2022)
  • ManyTypes4Py v0.7 Mir et al. (2022)
2021:
  • SySeVR Li et al. (2018a)
  • Ling&Wu et al. Ling et al. (2021)
  • Chen et al. Chen et al. (2021b)
  • MBPP, MathQA-Python Austin et al. (2021)
  • HumanEval Chen et al. (2021a)
  • APPS Hendrycks et al. (2021)
  • Berabi et al. Berabi et al. (2021)
  • CrossVul Nikitopoulos et al. (2021)
  • PYPIBUGS, RANDOMBUGS Allamanis et al. (2021)
  • D2A Zheng et al. (2021)
  • CodeQA Liu and Wan (2021)
  • Spider-DK Gan et al. (2021b)
  • KaggleDBQA Lee et al. (2021)
  • SEDE Hazoom et al. (2021)
  • Spider-Syn Gan et al. (2021a)
  • CoDesc Hasan et al. (2021)
  • Methods2Test Tufano et al. (2022)
  • Rozière et al. Rozière et al. (2022)
2020:
  • Lachaux&Roziere et al. Rozière et al. (2020)
  • $\mu$VulDeePecker Zou et al. (2020)
  • CosBench Yan et al. (2020)
  • PACS Heyman and Cutsem (2020)
  • Criteria2SQL Yu et al. (2020)
  • SQUALL Shi et al. (2020)
  • Hu et al. Hu et al. (2019)
  • CodeSearchNet Challenge Husain et al. (2019)
  • MIMICSQL Wang et al. (2020)
  • Atlas Watson et al. (2020)
  • Liu et al. Liu et al. (2022)
  • Android Agarwal et al. (2020)
  • CCSD Liu et al. (2021)
2019:
  • BFP Tufano et al. (2019)
  • SARD Lin et al. (2019)
  • Spider Yu et al. (2018)
  • JuICe Agashe et al. (2019)
  • Nguyen et al. Nguyen et al. (2019)
  • Lin et al. Lin et al. (2021)
  • Zhou et al. Zhou et al. (2019)
  • CoSQL Yu et al. (2019a)
  • SParC Yu et al. (2019b)
  • Malik Malik et al. (2019)
  • LeClair LeClair et al. (2019)
2018:
  • CoNaLa Yin et al. (2018)
  • DeepCom Hu et al. (2018a)
  • TL-CodeSum Hu et al. (2018b)
  • code-summarization-public Wan et al. (2018)
  • Russell et al. Russell et al. (2018)
  • VulDeePecker Li et al. (2018b)
  • Lin et al. Lin et al. (2018)
  • StaQC Yao et al. (2018)
  • Advising Finegan-Dollak et al. (2018)
  • ConCode Iyer et al. (2018)
  • NNGen Liu et al. (2018)
  • Gu et al. Gu et al. (2018)
2017:
  • QuixBugs Lin et al. (2017)
  • the DeepFix dataset Gupta et al. (2017)
  • Barone et al. Barone and Sennrich (2017)
2016:
  • CODE-NN Iyer et al. (2016)
  • Mou et al. Mou et al. (2016)
2015:
  • MANYBUGS, INTROCLASS Le Goues et al. (2015)
2014:
  • Defects4j Just et al. (2014)
  • BigCloneBench Svajlenko et al. (2014)
Appendix E Guideline
Finally, for ease of printing and use, we organized the guideline How2Bench into a clear, color-coded checklist (4 pages in total) that is easy to print, attached at the end of the paper.
See pages 1- of Figures/How-to-Bench.pdf

### v3

@@
-How Should I Build A Benchmark? Revisiting Code-Related Benchmarks For LLMs
+How Should We Build A Benchmark? Revisiting 274 Code-Related Benchmarks For LLMs
 Abstract
@@
 We propose How2Bench comprising a 55-criteria checklist as a set of guidelines to comprehensively govern the development of code-related benchmarks.
-Using How2Bench, we profiled 274 benchmarks released within the past decade and found concerning issues.
+Using How2Bench, we profiled 274 code-related benchmarks released within the past decade and found concerning issues.
 Nearly 70% of the benchmarks did not take measures for data quality assurance; over 10% did not even open source or only partially open source.
@@
 MyFramelinecolor=black, outerlinewidth=0.3pt, roundcorner=5pt, skipabove = 5.5pt, skipbelow = 5.5pt, innertopmargin=5.5pt, innerbottommargin=5.5pt, innerrightmargin=5.5pt, innerleftmargin=5.5pt, backgroundcolor=bg,
-How Should I Build A Benchmark?
-Revisiting Code-Related Benchmarks For LLMs
+How Should We Build A Benchmark?
+Revisiting 274 Code-Related Benchmarks For LLMs
 Jialun Cao1 , Yuk-Kit Chan2thanks:   Both authors contributed equally to this research.
@@
 We also observed these loopholes can propagate.
-Over 18% of the benchmarks serve as data sources (Figure 8) for subsequent benchmarks.
+Over 18% of the benchmarks serve as data sources for subsequent benchmarks (Figure 8).
 Therefore, the data quality of benchmarks affects their credibility and likely impacts future benchmarks.
@@
 (2024); Yuan et al.
-(2024, 2023a).
+(2024b, 2023a).
 A survey Chang et al.
@@
 Code-related benchmark development comprises five typical phases (Phase 0 - 4), as shown in Figure 1, explained in detail as follows.
-Phase 0. Design. At the beginning of benchmark development, it is vital to identify the motivation, the scope and the capabilities required by the application scenario of interest.
+Phase 0.
+Design.
+At the beginning of benchmark development, it is vital to identify the motivation, the scope and the capabilities required by the application scenario of interest.
+To achieve this objective, one needs to carefully consider the application scenarios, making sure these scenarios align with real-world demands.
+Also, it is also necessary to assess whether other benchmarks already exist that address similar tasks, and to identify any shortcomings they may possess.
+Furthermore, this new benchmark should be designed to evaluate specific LLMs’ capabilities; the crafted tasks are expected to reflect these capabilities.
 Phase 1.
@@
 3.2 Study Design
-Our study consists of four steps (Figure 2).
+Our study consists of four steps (Figure 2). All steps are explained as follows.
 Step 1.
@@
 This step is followed by profiling each selected benchmark through a thorough review of corresponding papers and examination of the released artifacts or homepages associated with these benchmarks.
-The phase is completed by reporting statistics that highlight overall trends, advantages, and disadvantages identified during the profiling, providing a clear and structured overview of existing benchmarks.
+The phase is completed by reporting statistics that highlight overall trends, pros, and cons identified during the profiling, providing a structured overview of existing benchmarks.
 Step 3.
@@
 The final step is a human study that evaluates the importance and practicality of How2Bench.
-This involves designing a questionnaire to gather diverse, logical insights, which is then distributed to a targeted audience.
+This involves designing a questionnaire by first initiating and iterating to gather diverse, logical insights, which is then distributed to a targeted audience.
 After collecting and filtering responses for quality, the data is analyzed to derive insights.
@@
 Key Statistics – According to our statistics among 270+ benchmarks, apparent research bias can be observed in terms of coding tasks, programming languages, and code granularities are observed (Appendix A.1).
-For example, as shown in Figure 11, 36.13% (99/274) are code generation benchmarks, followed by program repair 9.85% (27/274).
+For example, 36.13% (99/274) are code generation benchmarks, followed by program repair, with 9.85% (27/274).
 Also, during the focused case study (listed in Appendix C), we identified that 10% benchmarks have not explicitly specified the capabilities (e.g., intention understanding, program synthesis) to be evaluated, and 30% have not specified application scenarios the benchmark targets.
@@
 Lastly, Figure 12 shows that 58% (158/274) code-related benchmarks involve Python, followed by 39% (107/274) involving Java.
-In comparison, 31 programming languages are only covered by one benchmark, and less than five benchmarks cover other 19 programming languages.
+Yet, 31 programming languages are only covered by one benchmark, and less than five benchmarks cover other 19 programming languages.
 This observation consolidates the observation from previous works Cao et al.
@@
 Output validation methods and evaluation metrics must be carefully designed and reviewed to ensure they effectively measure the benchmark’s goals.
-Lastly, it suggests considering additional evaluation perspectives, such as safety and security checks, ensuring the code does not contain sensitive information.
+Lastly, it suggests considering additional evaluation perspectives, such as safety Wei et al.
+(2024); Yuan et al.
+(2024a) checks, ensuring the code does not contain sensitive information.
 Key Statistics – According to our statistics (Appendix A.3), the 270+ benchmarks exhibit numerous irregularities in their implementation, which could significantly threaten the reliability of the benchmarks.
@@
 (2023a), inadequate test coverage can lead to inflated evaluation results.
-However, according to our statistics, we observed that only 8.7% of benchmarks have considered test coverage when using test cases as oracles.
-It severely affects the reliability of findings on these benchmarks, potentially misguiding future research and application decisions based on these flawed assessments.
+However, we observed that only 8.7% of benchmarks have considered test coverage when using test cases as oracles (Appendix A.3).
+It severely affects the reliability of findings on these benchmarks, potentially misguiding future research and applications based on these flawed assessments.
 {mdframed}
-[style=MyFrame] ▲ Severity – Most benchmarks display severe loopholes in data preparation and curation processes.
-Essential steps such as data cleaning, process oversight, and quality checks are often neglected.
+[style=MyFrame] ▲ Severity – Most benchmarks display severe loopholes in data preparation and curation. Quality checks are often neglected.
 4.3 Guideline for Evaluation
@@
 Furthermore, because of the nondeterministic nature of LLMs, experiments should be repeated, and randomization strategies should be used to mitigate the effects of randomness and parameter configuration biases.
-Lastly, meticulously documented logs of the experimental process are advised to facilitate transparency and reproducibility, detailing everything from parameter settings to the specific LLM frameworks used.
+Lastly, meticulously documented logs of the experimental process are advised to facilitate transparency and reproducibility, detailing everything from parameter settings to the specific LLM pipelines such as vLLM (Kwon et al., 2023) used.
 Key Statistics – Among the 274 benchmarks, 183 of them are evaluated over LLMs.
@@
 Similarly, though 94.9% benchmarks were evaluated in a zero-shot manner, only 21.2% benchmarks were evaluated under few-shot, 8.8% under Chain-of-Thought and 2.6% under RAG (Appendix A.4).
+However, as shown in Figure 34, 73.3% representative benchmarks (Appendix C) do not validate whether the prompt they used is well-designed.
 Regarding the evaluation process, our statistics exposed that only 35.4% of benchmark evaluations have been repeated (Appendix A.4).
@@
 Figure 6: Guideline for Evaluation Analysis
-Explanation – We listed 10 criteria for the evaluation analysis phase, as shown in Figure 6.
+Explanation – The analysis of the experiment results is expected to be objective and comprehensive, hopefully providing insights or actionable advice.
+So, we listed 10 criteria for the evaluation analysis phase, as shown in Figure 6.
 Regarding the perspectives of analysis, inspired by classic measurement theory Suppes et al.
-(1962), we suggest four essential perspectives, including difficulty (whether a benchmark is appropriately challenging for LLMs), stability (whether the results are consistent through repeated trials), differentiability (whether benchmarks can differentiate the strengths and weaknesses of various LLMs).
+(1962), we suggest four essential perspectives, including difficulty (whether a benchmark is appropriately challenging for LLMs), stability (whether the results are consistent through repeated trials), differentiability (whether benchmarks can differentiate the strengths and weaknesses of various LLMs), and inspiration (e.g., the correlations between the upper-/down-stream coding tasks and LLM scores).
 Moreover, effective presentation of results using clear visual and textual descriptions is recommended to ensure the findings are understandable and actionable.
 The phase concludes with the suggestion to interpret and explain the results comprehensively, providing a basis for future research and application enhancements.
-Key Statistics – Because experimental analysis is relatively subjective and cannot be obtained through mechanical scanning, we focus on 30 representative focus benchmarks (Appendix C), covering the highest cited and latest benchmarks in the five tasks (Section 3.2 Step 3).
+Key Statistics – Because experimental analysis is relatively subjective and cannot be obtained through mechanical scanning, we focus on 30 representative focus benchmarks (Appendix C), covering the highest cited and latest benchmarks in top five tasks.
 Figure 37 shows an example from CruxEval Gu et al.
@@
 [style=MyFrame] ▲ Severity – The analysis of experimental data and the clarity of data presentation may receive less attention and worth consideration.
+Even in papers cited 1k+ times like MBPP Austin et al.
+(2021), there are instances of unclear evaluation analysis and display.
 4.5 Guideline for Benchmark Release
@@
 Furthermore, prompts, which are necessary for reproducibility, are not disclosed in 52.6% of the benchmarks (Figure 39).
-Not to mention the lack of public information on experimental settings and experimental parameters (Figure 43).
-Furthermore, only 16.7% of the benchmarks make their logged experimental results publicly available.
-See Appendix A.6 for details.
+Not to mention the lack of public information on experimental settings (Figure 32 and Figure 31) and experimental parameters (Figure 43).
+What is worse, 19.3% benchmarks do not setup licenses (Figure 44).
+The absence of licensing may lead to severe legal and ethical issues, potentially resulting in unauthorized use and distribution of proprietary technologies.
+Additionally, only 16.7% of the benchmarks make their logged experimental results publicly available (Appendix A.6).
 {mdframed}
@@
 A benchmark with contaminated data may yield overclaimed results, misleading the understanding of the LLMs’ capabilities.
-According to our statistics (Figure 23 on benchmarks from the year 2023 to 2024 (the duration when most LLMs were launched), most (78 %) benchmarks were not aware of and have not taken any measures to alleviate data contamination, being vulnerable to data contamination threat.
+According to our statistics (Figure 23 on benchmarks from the year 2023 to 2024 (the duration when most LLMs were launched), most (81.8 %) benchmarks were not aware of and have not taken any measures to alleviate data contamination, being vulnerable to data contamination threat.
+0
 Figure 23: Benchmark Distribution over Quality Assurance on Data Contamination
@@
 We summarize the number of LLMs that have been evaluated in each benchmark evaluation.
-As shown in Figure 29, the majority of benchmarks are tested with a smaller number of LLMs, as evidenced by the bars for 1 to 9 LLMs.
-Interestingly, there is a significant peak at the 10-49 LLM range, with 56 benchmarks falling into this category, suggesting a common practice of testing benchmarks extensively but not excessively.
-The drop in benchmarks tested with 50-99 and 100+ LLMs indicates that very extensive testing across numerous models is less common, potentially due to resource constraints or diminishing returns on insights gained from additional models.
+Among the 274 benchmarks, 183 of them are evaluated over LLMs, so we show the statistics over them.
+As shown in Figure 29, over 34% of the benchmarks were evaluated on fewer than 3 LLMs, with 11.48% benchmarks only evaluated on one LLM.
+Such evaluation results can hardly be generalized to other LLMs.
+Furthermore, more than half of the benchmarks studied fewer than 6 LLMs (51% = (21 +22 + 20 + 4 + 12+15)/183).
 Figure 29: Benchmark Distribution over LLM Experimented
@@
 Results Collection The responses were automatically collected through the online platform used for distribution.
-Survey Screening Since the requirement was for participants who have published papers, responses from those selecting "No" to having published a paper were excluded.
-Additionally, incomplete surveys where not all questions were answered were also considered invalid and excluded from the analysis.
+Survey Screening Since the requirement was for participants who have published papers, responses from those selecting “No” to having published a paper were excluded.
+Also, incomplete surveys where not all questions were answered were also considered invalid and excluded from the analysis.
 B.4 Interview Result Analysis
@@
 A breakdown of the respondents’ demographics is shown in Figure 50.
-And the detailed response for all 55 criteria in How2Bench are shown in Figure 51 and Figure 52.
+The detailed responses for all 55 criteria in How2Bench are shown in Figure 51 and Figure 52.
 Figure 49: Geographical Distribution of Interviewees